OpenStudy (anonymous):

HELP ME AND ILL WILL FAN AND MEDAL Solve. w – 8 = 32 A. 4 B. 24 C. 40 D. 256

OpenStudy (perl):

w = 8 + 32

OpenStudy (anonymous):

THX

OpenStudy (perl):

:)

OpenStudy (anonymous):

DO YOU LIKE MY CAR

OpenStudy (perl):

you own that?

OpenStudy (anonymous):

YES

OpenStudy (perl):

im not really a car person. what kind of car is it?

OpenStudy (anonymous):

BMW I8

OpenStudy (perl):

how much does it cost

OpenStudy (anonymous):

$135,700

OpenStudy (perl):

you dont own this car, lol

OpenStudy (anonymous):

Solve. k + 4.9 = 12.28 A. 7.38 B. 8.38 C. 16.18 D. 17.18

OpenStudy (anonymous):

OK LOL

OpenStudy (perl):

k = 12.28 - 4.9

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Which equation can be used to find the answer? A playground with four sides has a perimeter of 52 ft. Three of the sides have lengths of 9 ft, 16 ft, and 19 ft. What is the length of the fourth side? A. s – 9 + 16 + 19 = 52 s =26 26 ft B. s + 9 + 16 + 19 = 52 s = 8 8 ft C. s – 52 = 9 + 16 – 19 s = 58 58 ft D. s – 9 – 16 – 19 = 52 s = 96 96 ft

OpenStudy (perl):

fan and medal me, make new question. you cant copy and paste special type

OpenStudy (perl):

cya , make a new question

OpenStudy (anonymous):

OK